OpenStudy (anonymous):

A school has 70 sixth graders, 65 seventh graders, and 75 eighth graders. In each grade, 20% of the students live within 1 mile of school. What is the total number of students that live more than 1 mile from school? A 42 B 168 C 190 D 210

OpenStudy (princeharryyy):

80percent of(70 + 65 + 75) =>(4/5)*(210) = 4*42 = 168 @lilprincess